A water sample has 5 mg/L of Nitrate, 10 mg/L of Potassium, 60 mg/L of Calcium, and 60 mg/L of Magnesium. If the alkalinity is 394 mg/L of CaCO3, calculate the Total Hardness, Carbonate Hardness, and Noncarbonate Hardness using all polyvalent cations.
Answer #1

In the problem it is given that the concentration of

Pottasium (K+) - 10 mg/l

Calcium (Ca+) - 60 mg/l

Magnisium (Mg+) - 60 mg/l

Total Alkalanity ( TA) - 394 mg/l

Total Hardness (TH)= Ca2+ ( Eq.wt of CaCO3) / Eq. wt of Ca2+ + Mg+( Eq.wt of CaCO3) / Eq. wt of Mg2+ +

K+( Eq.wt of CaCO3) / Eq. wt of K+

  Eq.wt. of CaCO3 =50 , Eq.wt. of Ca2+   = 20 , Eq.wt. of Mg2+   = 12 , Eq.wt. of K+   = 39

So Total Hardness (TH) = 60 *50 /20 + 60 *50/12 + 10 *50/39 = 413 mg/l

Carbonate Hardness (CH) = TA or TH whichever is less , here TA TH

So Carbonate Hardness = 394 mg/l

Non Carbonate Hardness = TH - TA = 413 - 394 6mg/l = 19 mg/l

   

Total Hardness 413 mg/l
Carbonate Hardness 394 mg/l
Non Carbonate Hardness 19 mg/l